OpenStudy (anonymous):

rationalize square root of 5 over square root of 11

OpenStudy (anonymous):

\[\sqrt{55}/11\]

OpenStudy (anonymous):

\[\sqrt{5}/\sqrt{11}=\frac{\sqrt{5}*\sqrt{11}}{\sqrt{11}*\sqrt{11}}=\sqrt{55}/11\]